OpenStudy (anonymous):

The sum of three consecutive integers is 228. Find the integers

OpenStudy (anonymous):

x+x+1+x+2=228

OpenStudy (anonymous):

x=1st integer x+1 = 2nd integer x+2 = 3rd integer

OpenStudy (anonymous):

consecutive integers are basically spelt x+x+1+x+2. In your case, it would all equal to 228, and the three integers are x, x+1, and x+2. so your equation would be: x+x+1+x+2=228. do you know how to solve for x?

OpenStudy (anonymous):

so, you could simplify the equation, by reducing it until the equation looked something like this: 3x+3=228. What I did here was just combining LIKE TERMS, are you familiar with that? Then, you solve for x. subtract the three from 228, in order to get x alone. your result will turn up as : 3x=225 Note that 3x is 3 MULTIPLIED with x. In order to get x alone, what do you do? You divide both sides (both sides of the equation) by 3. Your result should be: x=75. You are not done yet. Remember that your consecutive integers are x, x+1, and x+2. Now, you just plug it in. Your first integer,x, will be 75, the second, 76 (x+1), and lastly, 77 (x+2). I hope I helped. If you have any questions, feel free to ask. I'll be happy to help! (;